A body of mass 0.120 kg and a velocity of 0.50 m/s hits another identical body at rest, and the two recoils forming 90 degrees between them. If the final velocity of one of the balls are directed 30 degrees above horizontal, find their speeds after collision.



1
Expert's answer
2022-01-10T09:09:59-0500

by momentum conservation:

"m_1v_1=m_1v_1'+m_2v_2"

where m1, m2 are masses,

v1, v1' are speeds of first body before and after collision

v2 is speed of second body after collision


then:

for x-direction:

"v_1'cos30\degree +v_2cos60\degree=v_1=0.5"

for y-direction:

"v_1'sin30\degree =v_2sin60\degree"


"v_1'\sqrt 3+v_2=1"

"v_1'=v_2\sqrt 3"


speeds after collision:

"v_2=1\/4=0.25" m/s

"v_1'=0.25\sqrt 3=0.43" m/s